### v1

@@ Abstract
The (strong) isometric path complexity is a recently introduced graph invariant that captures how arbitrary isometric paths (i.e. shortest paths) of a graph can be viewed as a union of few “rooted" isometric paths (i.e. isometric paths with a common end-vertex). It is known that this parameter can be computed optimally in polynomial time. Seemingly unrelated graph classes studied in metric graph theory (e.g. hyperbolic graphs), geometric intersection graph theory (e.g. outerstring graphs), and structural graph theory (e.g. (theta, prism, pyramid)-free graphs) have been shown to have bounded strong isometric path complexity [Chakraborty et al., MFCS ’23].
We show that important graph classes studied in coarse graph theory (as introduced by [Georgakopoulos & Papasoglu ’23]) have bounded strong isometric path complexity. Specifically, we show that the strong isometric path complexity of $K_{2,t}$-asymptotic minor-free graphs is bounded. In fact, we prove a more general result. Let $U_{t}$ denote the graph obtained by adding a universal vertex to a path of $t-1$ edges. We show that the strong isometric path complexity of $U_{t}$-asymptotic minor-free graphs is bounded. This implies that $K_{4}^{-}$-asymptotic minor-free graphs, i.e. graphs that are quasi-isometric to a cactus [Fujiwara & Papasoglu ’23] have bounded strong isometric path complexity. On the other hand, $K_{4}$-minor-free graphs have unbounded strong isometric path complexity.
We also show that graphs whose all induced cycles of length at least 4 have the same length (also known as monoholed graphs as defined by [Cook et al., JCTB ’24]) form a subclass of $U_{4}$-asymptotic minor-free graphs. Hence, the strong isometric path complexity of monoholed graphs is bounded. On the other hand, we show that even-hole free graphs have unbounded strong isometric path complexity.
We also investigate which graph operations preserve strong isometric path complexity. We show that the strong isometric path complexity is preserved under the fixed power and line graph operators, two important graph operations. We also show that the clique-sums of finitely many graphs with strong isometric path complexity at most $k$, yields a graph with strong isometric path complexity at most $183k+$ 18.
@@ 1 Introduction
The study of the distance-metric of graphs is at the heart of metric graph theory [7] and finds numerous practical applications. To captures how arbitrary isometric paths (i.e. shortest paths) of an unweighted graph can be viewed as a union of few “rooted" isometric paths (i.e. isometric paths with a common end-vertex), a graph invariant known as isometric path complexity was introduced in [10]. Formally, it is defined as follows.
Given an unweighted and undirected graph $G$ and a vertex $v\in V(G)$, an isometric path $P$ of $G$ is $v$-rooted if $v$ is one of the end-vertices of $P$. The isometric path complexity of a connected graph $G$, denoted by $ipco\left(G\right)$, is the minimum integer $k$ such that there exists a vertex $v\in V(G)$ satisfying the following property: the vertices of any isometric path of $G$ can be covered by $k$ many $v$-rooted isometric paths.
A stronger variant of the isometric path complexity which is the main topic of this paper is also defined in [10] as follows.
@@
Clearly, for any graph $G$, $ipco\left(G\right)\leq sipco\left(G\right)$. We say that the isometric path complexity (resp. strong isometric path complexity) of a graph class $\mathcal{G}$ is bounded if there is a constant $k$ that depends only on $\mathcal{G}$ such that for all $G\in\mathcal{G}$, the isometric path complexity (resp. strong isometric path complexity) is at most $k$.
@@
Motivation: Originally, these parameters were conceived to propose approximation algorithms for the problem Isometric Path Cover, (also known as Shortest Path Cover) where the objective is to find a minimum-cardinality set of isometric paths so that each vertex of the graph belongs to at least one of the solution paths. Isometric Path Cover has been studied recently in the algorithmic graph theory community [11, 10, 12] and finds applications in machine learning [32], cop decompositions (in relation with the celebrated Cops and Robber game [1]), bus route design [11] etc. It was shown that Isometric Path Cover admit a polynomial-time $c$-approximation algorithm on graphs with isometric path complexity at most $c$ [10] (while being NP-hard even on chordal graphs). Unlike some other metric graph parameters (e.g. treelength), the (strong) isometric path complexity can be computed optimally in polynomial time [10].
@@
From the structural perspective, it turns out that many important graph classes such as hyperbolic graphs, outerstring graphs, and (theta,prism,pyramid)-free graphs have bounded strong isometric path complexity [10]. These graph classes are incomparable to each other with respect to inclusion, and isometric path complexity being bounded is perhaps the first known interesting property shared by these seemingly unrelated graph classes. Hyperbolic graphs are important in metric graph theory [7, 26] and include classic graph classes such as chordal graphs, and more generally, graphs with bounded tree-length). But hyperbolic graphs exclude arbitrarily large cycles as isometric subgraphs. Outerstring graphs (and its subclasses) are studied in geometric intersection graph theory and computational geometry [6, 8, 9, 31]. The class of (theta,prism,pyramid)-free graphs has been studied in structural graph theory [2, 15]. Note that (as opposed to hyperbolic graphs) outerstring graphs and (theta,prism,pyramid)-free graphs may contain arbitrarily many large isometric cycles, although in a structured manner. See Figure 1 for graph classes with bounded strong isometric path complexity. In this paper, we show that the strong isometric path complexity is also important in coarse graph theory.
@@
Figure 1: Inclusion diagram for graph classes. If a class $A$ has an upward path to class $B$, then $A$ is included in $B$. Constant bounds for the strong isometric path complexity on graph classes and graph operations marked with * are contributions of this paper. The boxes marked in gray indicate graph classes with unbounded isometric path complexity. Elliptical boxes represent graph operations.
@@
Asymptotic minors: Recently, Georgakopoulos and Papasoglu [25] gave an overview of the area called coarse graph theory, where the objective is to “look at graphs from far away” and to understand their “large-scale geometry” [25]. Many problems and questions posed in [25] involve asymptotic minors and fat minors, and have garnered recent attention [24, 3, 5, 4, 20]. A graph $G$ contains a graph $H$ as a $K$-fat minor if it is possible to find in $G$, (i) a collection of connected subgraphs called branch sets corresponding to $V(H)$, and a collection of connected sets called branch paths corresponding to $E(H)$ such that (i) after contracting each branch set to a vertex, and each branch path to an edge, we obtain a copy of $H$, and (ii) the above sets are at pairwise distance at least $K$, except for incident branch set–branch paths pairs, which must be at distance 00 by definition. A graph class $\cal G$ contains $H$ as an asymptotic minor if for every integer $K$, there exists a graph $G\in\cal G$, such that $G$ contains $H$ as a $K$-fat minor. Otherwise, $\cal G$ is $H$-asymptotic minor-free.
It was conjectured in [25] that $H$-asymptotic minor-free graphs are quasi-isometric to $H$-minor-free graphs for every graph $H$. (A quasi-isometric mapping of a graph $G$ to a graph $H$ is a mapping of $V(G)$ to $V(H)$ that preserves the distance up to a constant multiplicative and additive distortion.) If true, coarse analogues of various results from graph minor theory would be true as well. However, the above conjecture has been proven to be false in general [20]. On the other hand, the above conjecture is true for $H=K_{4}$ [4], $H=K_{4}^{-}$ [24, 4], and $H=K_{2,3}$ [14].
@@
In this paper, we consider a super class of $K_{2,3}$-asymptotic minor-free graphs and $K_{4}^{-}$-asymptotic minor free graphs and show that its strong isometric path complexity is bounded. Let $U_{t}$ denote the graph obtained by taking a path of length $t$ and introducing a universal vertex. See Figure 2 for an example. We prove the following.
Figure 2: The graph $U_{5}$.
@@
We note that $K_{2,t}$-asymptotic minor-free graphs are also $U_{t}$-asymptotic minor-free graphs. (See Proposition 1 for a proof.) Hence, Theorem 1 implies the following.
@@
The class of $K_{2,3}$-asymptotic minor-free graphs and its subclasses (e.g. $K_{2,3}$-induced minor-free graphs, universally-signable graphs, outerstring graphs) have been studied in the literature [14, 9, 6, 17, 29, 19]. We note that $K_{2,t}$-asymptotic minor-free graphs are strictly included in the class of $U_{t}$-asymptotic minor-free graphs. For an integer $k\geq 2$, let $F_{k}$ denote the graph obtained by taking $k$ copies of paths of length $k$ and identifying the end-vertices of the paths. Let $\mathcal{F}=\displaystyle\bigcup\limits_{k\geq 2}F_{k}$. Clearly, $\cal F$ is $U_{t}$-asymptotic minor-free but contains $K_{2,t}$ as an asymptotic minor for all $t\geq 1$. observe that all graphs in $\cal F$ are monoholed i.e. the lengths of all induced cycles in the graph are the same.
Monoholed graphs and even-hole free graphs: A hole is an induced cycle of length at least 4. The study of graphs whose hole lengths are restricted has been pursued in Structural Graph Theory [18, 27, 35]. A graph $G$ is $\ell$-holed if all holes of $G$ are of length $\ell$. A graph $G$ is monoholed if it is $\ell$-holed for some integer $\ell$. For any odd integer $\ell$, $\ell$-holed graphs are also even-hole free graphs (i.e. graphs with no holes of even length) and, for even integers $\ell$, it is known that $\ell$-holed graphs are perfect [27], that is, do not contain any odd holes or complements of an odd hole. Studies on even-hole free graphs [33] and perfect graphs [16] have yielded many powerful decomposition theorems and algorithmic techniques for fundamental problems such as Coloring, Independent Set, and Clique.
@@
The above theorem implies that Isometric Path Cover admits a constant factor approximation algorithm on monoholed graphs. Some other algorithmic results on monoholed graphs are known. Maximum Independent Set can be solved in polynomial time on monoholed graphs [18]. Clique can be solved in polynomial time for $\ell$-holed graphs with $\ell\geq 5$, as there are only polynomially many maximal cliques in such graphs. The computational complexity of Coloring is open on these graphs. Already, optimally coloring the special case of rings is a difficult (but solved) problem [28].
Next, we consider two superclasses of monoholed graphs: perfect graphs and even-hole free graphs. Bipartite graphs (which form a subset of both these classes) have unbounded isometric path complexity. Indeed, it can be easily checked that the isometric path complexity of an $(n\times n)$-grid is $\Omega(n)$. Hence, perfect graphs have unbounded isometric path complexity. We show that the isometric path complexity of even-hole free graphs is also unbounded.
@@
The class of even-hole free graphs with maximum degree at most 4 has unbounded isometric path complexity.
Graph operations: Quasi-isometry is a generalisation of biLipschitz maps that allows additive distortion. Embeddings with multiplicative (i.e. biLipschitz embeddings) or additive distortion are an area of active research [25, 24, 3, 5, 23, 21, 13]. Hyperbolic graphs are closed under quasi-isometry [26]. For a fixed graph $H$, $H$-asymptotic minor-free graphs are also closed under quasi-isometry [25]. Therefore, we investigate whether the class of graphs with bounded (strong) isometric path complexity is closed under quasi-isometry.
We prove that two popular quasi-isometric graph operations: graph powers and line graphspreserve strong isometric path complexity. For an integer $r\geq 1$ and a graph $G$, the $r^{th}$ power of $G$, denoted as $G^{r}$, is the graph with vertex set $V(G)$ where two vertices are adjacent in $G^{r}$ if they are at distance at most $r$ in $G$. The line graph of $G$, denoted as $L\left(G\right)$, has vertex set $E(G)$ and two vertices are adjacent in $L\left(G\right)$ if the corresponding edges share an end-vertex in $G$. Clearly, for a fixed integer $r$, any graph $G$ and its $r^{th}$ power ${G}^{r}$ are quasi-isometric. Similarly, any graph $G$ and its line graph $L\left(G\right)$ are also quasi-isometric. Line graphs and fixed powers of graphs are important in various areas of graph theory including metric graph theory [7]. We prove the following.
@@
    the isometric path complexity (resp. strong isometric path complexity) of $L\left(G\right)$ is at most $23k+$ 2. Moreover, for every $t\geq 1$ there exists a graph $H_{t}$ with $ipco\left(H_{t}\right)=t$ and $ipco\left(L\left(H_{t}\right)\right)=2t-1$.
@@
If two graphs $G$ and $H$ each contain cliques of equal size, the clique-sum of $G$ and $H$ is done from the disjoint union of $G$ and $H$ by identifying pairs of vertices in these two cliques to form a single shared clique. The resulting graph is $G\oplus H$. Clique-sums are important in graph minor theory. This operation is also studied in the context of decomposition theorems (e.g. for chordal graphs, universally-signable graphs [17] etc.). The clique-sum of two hyperbolic graphs (resp. $H$-asymptotic minor-free graphs) yields a hyperbolic graphs (resp. $H$-asymptotic minor-free graphs). However, it is known that the isometric path complexity is not closed under clique-sum [10]. We prove the contrary for strong isometric path complexity. For graphs $G_{1},G_{2},\ldots,G_{t}$, denote by $\bigoplus\limits_{i=1}^{t}G_{i}$ the graph obtained by taking pairwise clique sums of $G_{1},G_{2},\ldots,G_{t}$. (Note that it is not necessary that all graphs are identified using a single clique.)
@@
Let $G_{1},G_{2},\ldots,G_{t}$ be graphs with $sipco\left(G_{i}\right)\leq k$ for all $i\in[t]$. Then the strong isometric path complexity of $G=\displaystyle\bigoplus\limits_{i=1}^{t}G_{i}$ is at most $183k+$ 18.
@@
Organization of the paper: In Section 2, we recall some definitions and some results. In Section 3 we prove Theorems 1, 2 and 3. In Section 4 we prove Theorems 4 and 5. We conclude in Section 5.

### v2

@@ Abstract
The (strong) isometric path complexity is a recently introduced graph invariant that captures how arbitrary isometric paths (i.e., shortest paths) of a graph can be viewed as a union of a few “rooted" isometric paths (i.e., isometric paths with a common end-vertex). It is known that this parameter can be computed optimally in polynomial time. Seemingly unrelated graph classes studied in metric graph theory (e.g. hyperbolic graphs), geometric intersection graph theory (e.g. outerstring graphs), and structural graph theory (e.g. (theta, prism, pyramid)-free graphs) have been shown to have bounded strong isometric path complexity [Chakraborty et al., Disc. Math. ’25], which renders this parameter promising in the realm of metric graph theory.
We show that important graph classes studied in coarse graph theory have bounded strong isometric path complexity. Specifically, we show that the strong isometric path complexity of $K_{2,t}$-asymptotic minor-free graphs is bounded. In fact, we prove a more general result. Let $U_{t}$ denote the graph obtained by adding a universal vertex to a path of t1t-1 edges. We show that the strong isometric path complexity of $U_{t}$-asymptotic minor-free graphs is bounded. This implies that $K_{4}^{-}$-asymptotic minor-free graphs, i.e., graphs that are quasi-isometric to a cactus [Fujiwara & Papasoglu ’23], have bounded strong isometric path complexity. On the other hand, $K_{4}$-minor-free graphs have unbounded strong isometric path complexity. Hence, for a graph $H$ on at most four vertices, $H$-asymptotic minor-free graphs have bounded (strong) isometric path complexity if and only if $H\not=K_{4}$.
We show that graphs whose all induced cycles of length at least 4 have the same length (also known as monoholed graphs as defined by [Cook et al., JCTB ’24]) form a subclass of $U_{4}$-asymptotic minor-free graphs. Hence, the strong isometric path complexity of monoholed graphs is bounded. On the other hand, we show that even-hole free graphs have unbounded strong isometric path complexity.
We investigate which graph operations preserve strong isometric path complexity. We show that the strong isometric path complexity is preserved under the fixed power and line graph operators, two important graph operations. We also show that the clique-sums of finitely many graphs with strong isometric path complexity at most $k$, yield a graph with strong isometric path complexity at most $3k+18$.
@@ 1 Introduction
All graphs considered in this paper are simple, undirected, unweighted, and finite. The study of the distance-metric of graphs is at the heart of metric graph theory [9] and finds numerous practical applications in computer science [9]. To capture how arbitrary isometric paths (i.e., shortest paths) of an unweighted graph can be viewed as a union of few “rooted” isometric paths (i.e., isometric paths with a common end-vertex), a graph invariant known as isometric path complexity was introduced in [17], where it was shown to have interesting properties and promising applicability. Formally, it is defined as follows.
Given a graph $G$ and a vertex $v\in V(G)$, an isometric path $P$ of $G$ is $v$-rooted if $v$ is one of the end-vertices of $P$. The isometric path complexity of a connected graph $G$, denoted by $ipco\left(G\right)$, is the minimum integer $k$ such that there exists a vertex $v\in V(G)$ satisfying the following property: the vertices of any isometric path of $G$ can be covered by $k$ many $v$-rooted isometric paths.
A stronger variant of the isometric path complexity, which is the main topic of this paper, was defined in [17] as follows.
@@
We say that the isometric path complexity (resp. strong isometric path complexity) of a graph class $\mathcal{G}$ is bounded if there is a constant $k$ that depends only on $\mathcal{G}$ such that for all $G\in\mathcal{G}$, the isometric path complexity (resp. strong isometric path complexity) is at most $k$. Clearly, for any graph $G$, the isometric path complexity is upper bounded by the strong isometric path complexity. The gap between the isometric path complexity and the strong isometric path complexity can be arbitrarily large [17].
Goal of the paper. The (strong) isometric path complexity is a promising metric-based graph parameter, as it captures several interesting and well-studied graph classes (that seemingly did not have much in common) under the same umbrella. Moreover, they can be computed in polynomial time and have interesting algorithmic applications. (See the next paragraph for details.) Moreover, these concepts find relevance in the recently introduced coarse graph theory. Hence, the properties of these parameters deserve to be studied further. In particular, the frontier between graph classes with bounded and unbounded values of the parameter, as well as the behavior of standard graph operations with respect to the parameters, are of interest. Our goal is thus to conduct such studies, with a particular focus on metric-based graph classes and graph operations, in order to extend the applicability of these concepts.
@@
Related work. Originally, the notion of isometric path complexity was conceived to propose approximation algorithms for the problem Isometric Path Cover [14], (also known as Shortest Path Cover [32]) where the objective is to find a minimum-cardinality set of isometric paths so that each vertex of the graph belongs to at least one of the solution paths. Isometric Path Cover has been studied recently in the algorithmic graph theory community [14, 17, 15] and finds applications in machine learning [45], cop decompositions (in relation with the celebrated Cops and Robber game [1]), bus route design [14] etc. It was shown that Isometric Path Cover admits a polynomial-time $c$-approximation algorithm on graphs with isometric path complexity at most $c$ [17] (while being NP-hard even on chordal graphs, which have isometric path complexity at most 4). Unlike some other metric graph parameters (e.g. treelength), the (strong) isometric path complexity can be computed optimally in polynomial time [17].
@@
From the structural perspective, it turns out that many important graph classes such as hyperbolic graphs, outerstring graphs, and (theta, prism, pyramid)-free graphs have bounded strong isometric path complexity [17]. These graph classes are important on their own and are incomparable to each other with respect to inclusion. To our knowledge, having bounded isometric path complexity is the first known interesting common feature shared by these seemingly unrelated graph classes.
The notion of $\delta$-hyperbolicity [37] measures how “treelike” a graph is from a metric point of view (we omit the formal definition in this paper). Graphs with constant $\delta$-hyperbolicity are called hyperbolic graphs. From a theoretical perspective, many popular graph classes like interval graphs, chordal graphs, $\alpha_{i}$-metric graphs [29], graphs with bounded tree-length [28], link graphs of simple polygons  [20] have constant $\delta$-hyperbolicity. From a practical perspective, the study of $\delta$-hyperbolicity of graphs is motivated by the fact that many real-world graphs are tree-like [3, 4, 39] or have small $\delta$-hyperbolicity [11, 31, 42]. However, hyperbolic graphs exclude arbitrarily large cycles as isometric subgraphs.
Outerstring graphs (and its subclasses) are studied in geometric intersection graph theory and computational geometry [8, 10, 13, 44]. Popular graph classes like outerplanar graphs and circle graphs are subclasses of outerstring graphs. The class of (theta, prism, pyramid)-free graphs has been studied in structural graph theory [2, 21] as a way to extend studies of hereditary graph classes, based on the vast literature around perfect graphs, cutsets, decomposition theorems, etc.
Note that (as opposed to hyperbolic graphs) outerstring graphs and (theta, prism, pyramid)-free graphs may contain arbitrarily many large isometric cycles, although in a structured manner. See Figure˜1 for graph classes with (un)bounded strong isometric path complexity. In this paper, we show that the strong isometric path complexity of graph classes important in coarse graph theory is bounded.
@@
Figure 1: Inclusion diagram for graph classes. If a class AA has an upward path to class $B$, then $A$ is included in $B$. Constant bounds for the strong isometric path complexity on graph classes and graph operations marked with * are contributions of this paper. The boxes marked in gray indicate graph classes with unbounded isometric path complexity. Elliptical boxes represent graph operations.
@@
Asymptotic minors and quasi-isometry. Recently, Georgakopoulos and Papasoglu [36] stated a number of results and conjectures about coarse properties of graphs, establishing a basis of a new area called coarse graph theory. In this area, the objective is to “look at graphs from far away” and to understand their “large-scale geometry”. Many problems and questions posed in [36] involve asymptotic minors and fat minors, and have garnered recent attention [35, 5, 7, 6, 26]. A graph $G$ contains a graph $H$ as a $K$-fat minor if it is possible to find in $G$, (i) a collection of connected subgraphs called branch sets corresponding to $V(H)$, and a collection of paths called branch paths corresponding to $E(H)$ such that (i) after contracting each branch set to a vertex, and each branch path to an edge, we obtain a copy of $H$, and (ii) the above sets are at pairwise distance at least $K$, except for incident branch set–branch paths pairs, which must be at distance 0 by definition. A graph class $\cal G$ contains $H$ as an asymptotic minor if for every positive integer $K$, there exists a graph $G\in\cal G$, such that $G$ contains $H$ as a $K$-fat minor. Otherwise, $\cal G$ is $H$-asymptotic minor-free. The structure of $H$-asymptotic minor-free graphs has been studied in particular when $H=K_{4}$ [6], $H=K_{4}^{-}$ [35, 6], and $H=K_{2,3}$ [19].
Quasi-isometry333 An $(M,A)$-quasi-isometry between graphs $G$ and $H$ is a map $f\colon V(G)\rightarrow V(H)$ such that the following holds for fixed constants $M\geq 1,A\geq 0$, (a) $M^{-1}\mathsf{d}\left(x,y\right)-A\leq\mathsf{d}\left(f(x),f(y)\right)\leq M\mathsf{d}\left(x,y\right)+A$for every $x,y\in V(G)$; and (b) for every $z\in V(H)$ there is $x\in V(G)$ such that $\mathsf{d}\left(z,f(x)\right)\leq A$. The graphs $G$ and $H$ are quasi-isometric if there exists a quasi-isometric map $f$ between $V(G)$ and $V(H)$. between graph classes is an important notion in coarse graph theory. Quasi-isometry is a generalisation of biLipschitz maps that allows additive distortion. Embeddings with multiplicative (i.e. biLipschitz embeddings) or additive distortion are an area of active research [36, 35, 5, 7, 34, 30, 18]. Hyperbolic graphs are closed under quasi-isometry [37]. For a fixed graph $H$, the property of excluding $H$ as an asymptotic minor is invariant under applying quasi-isometries [36]. On the contrary, the property of bounded (strong) isometric path complexity is not preserved under quasi-isometry (see Remark˜1). However, this notion has been used in some proofs regarding quasi-isometry between graph classes. For example, the notion of strong isometric path complexity has been used to show that given a $K_{2,3}$-induced minor-free [^4] A graph $G$ contains a graph $H$ as an induced minor, if $H$ can be obtained from $G$ by deleting vertices and contracting edges. Otherwise, $G$ is $H$-induced minor-free. graph $G$, it is possible to construct a graph $H$ in $O(nm)$-time such that $G$ is quasi-isometric to $H$ with a constant additive distortion and $H$ has treewidth at most two [16]. In the special case of universally signable graphs [47], the above quasi-isometry can be obtained in linear time. This further implies a truly-subquadratic additive approximation algorithm for Diameter [^5] Diameter is the algorithmic problem of finding the longest isometric path in a graph. on universally signable graphs. In contrast, assuming the Strong Exponential Time Hypothesis (SETH), the diameter of split graphs (a very restricted class of universally signable graphs), cannot be computed optimally in truly sub-quadratic time [12]. However, the fact that $K_{2,3}$-induced minor-free graphs (more generally $K_{2,3}$-asymptotic minor-free graphs) are quasi-isometric to a cactus with additive distortion can be alternatively derived using known results [35, 19].
@@
In this paper, we consider a superclass of $K_{2,3}$-asymptotic minor-free graphs and $K_{4}^{-}$-asymptotic minor-free graphs and show that its strong isometric path complexity is bounded. Let $U_{t}$ denote the graph obtained by taking a path with $t$ vertices and introducing a universal vertex. See Figure˜2(a) for an example. We prove the following.
$\vdots$
(a) (b)
Figure 2: (a) The graph $U_{5}$. (b) The graph $F_{5}$. The dashed edges indicate paths.
@@
Since (theta, prism, pyramid)-free graphs and outerstring graphs are $U_{3}$-asymptotic minor-free, the above theorem generalises two results of [17]. It is known that $K_{3}$-asymptotic minor-free graphs have bounded treelength [36, 28], and therefore have bounded strong isometric path complexity. The above theorem implies that $U_{3}$-asymptotic minor-free graphs have bounded strong isometric path complexity. On the other hand, it is known from [17] (and also implied from Theorem˜7) that $K_{4}$- minor-free graphs have unbounded (strong) isometric path complexity. Hence, we have the following corollary.
Corollary 3.
Let $H$ be a graph with at most four vertices. Then $H$-asymptotic minor-free graphs have bounded strong isometric path complexity if and only if $H\not=K_{4}$.
Using a result from [17], Theorem˜2 also implies the following corollary.
Corollary 4.
Isometric Path Cover admits a constant-factor approximation algorithm on $U_{t}$-asymptotic minor-free graphs for all $t\geq 1$.
We note that $K_{2,t}$-asymptotic minor-free graphs are also $U_{t}$-asymptotic minor-free graphs. See Proposition˜11 for a proof. Hence, Theorem˜2 implies the following.
@@
[…]
Monoholed graphs and even-hole free graphs. A hole is an induced cycle of length at least 4. The study of graphs whose hole lengths are restricted has been pursued in Structural Graph Theory [24, 38, 49]. A graph $G$ is $\ell$-holed if all holes of $G$ are of length $\ell$. A graph $G$ is monoholed if it is $\ell$-holed for some integer $\ell$. For any odd integer $\ell$, $\ell$-holed graphs are also even-hole free graphs (i.e. graphs with no holes of even length) and, for even integers $\ell$, it is known that $\ell$-holed graphs are perfect [38], that is, do not contain any odd holes or complements of an odd hole as induced subgraphs. Studies on even-hole free graphs [46] and perfect graphs [22] have yielded many powerful decomposition theorems and algorithmic techniques for fundamental problems such as Coloring, Independent Set, and Clique.
@@
The above theorem implies that Isometric Path Cover admits a constant-factor approximation algorithm on monoholed graphs. Some other algorithmic results on monoholed graphs are known. Maximum Independent Set can be solved in polynomial time on monoholed graphs [24]. Clique can be solved in polynomial time for $\ell$-holed graphs with $\ell\geq 5$, as there are only polynomially many maximal cliques in such graphs. The computational complexity of Coloring is open on these graphs. Already, optimally coloring the special case of rings (a special type of monoholed graphs that is important in their study) is a difficult (but solved) problem [40].
Next, we consider two superclasses of monoholed graphs: perfect graphs and even-hole free graphs. Bipartite graphs (which form a subset of both these classes) have unbounded isometric path complexity. Indeed, it can be easily checked that the isometric path complexity of an $(n\times n)$-grid is $\Omega(n)$. (Indeed, for each vertex $v$ of a $(n\times n)$-grid there exists an isometric path $P_{v}$ containing $\Omega(n)$ many vertices which are equidistant from $v$ and no two such vertices can be covered by one $v$-rooted isometric path.) Hence, perfect graphs have unbounded isometric path complexity. We show that the isometric path complexity of even-hole free graphs is also unbounded.
@@
The class of even-hole free graphs with maximum degree at most $3$ has unbounded isometric path complexity.
Graph operations. In this paper, we study the behaviour of (strong) isometric path complexity under popular graph operations. We prove that two popular quasi-isometric graph operations: graph powers and line graphs preserve the property of having bounded strong isometric path complexity. For an integer $r\geq 1$ and a graph $G$, the $r^{th}$ power of $G$, denoted as $G^{r}$, is the graph with vertex set $V(G)$ where two vertices are adjacent in $G^{r}$ if they are at distance at most $r$ in $G$. The line graph of $G$, denoted as $L\left(G\right)$, has vertex set $E(G)$ and two vertices are adjacent in $L\left(G\right)$ if the corresponding edges share an end-vertex in $G$. Clearly, for a fixed integer $r$, any graph $G$ and its $r^{th}$ power ${G}^{r}$ are quasi-isometric. Similarly, any graph $G$ and its line graph $L\left(G\right)$ are also quasi-isometric. Line graphs and fixed powers of graphs are important in various areas of graph theory including metric graph theory [9]. We prove the following.
@@
    the isometric path complexity (resp. strong isometric path complexity) of $L\left(G\right)$ is at most $3k+1$. Moreover, for every $t\geq 1$ there exists a graph $H_{t}$ with $ipco\left(H_{t}\right)=t$ and $ipco\left(L\left(H_{t}\right)\right)=2t-1$.
@@
If two graphs $G$ and $H$ each contain cliques $C_{G}$ and $C_{H}$, respectively, of equal size, the clique-sum of $G$ and $H$ is obtained from the disjoint union of $G$ and $H$ by identifying pairs of vertices of $C_{G}$ and $C_{H}$ to form a single shared clique. Note that for different $C_{G}$ and $C_{H}$, the resulting graph can be different. Let $G\oplus H$ denote a graph that can be obtained by clique sum of $G$ and $H$. Clique-sums are important in diverse areas of graph theory, e.g., Wagner’s Theorem on characterization of $K_{5}$-minor-free graphs [48], graph minor theory [27] etc. This operation is also studied in the context of decomposition theorems (e.g. for chordal graphs, universally-signable graphs [23] etc.). The clique-sum of two hyperbolic graphs (resp. $H$-asymptotic minor-free graphs) yields a hyperbolic graph (resp. $H$-asymptotic minor-free graphs). We prove a similar result for graphs with bounded strong isometric path complexity. For graphs $G_{1},G_{2},\ldots,G_{t}$, denote by $\bigoplus\limits_{i=1}^{t}G_{i}$ the graph obtained by taking pairwise clique sums of $G_{1},G_{2},\ldots,G_{t}$. (Note that it is not necessary that all graphs are identified using a single clique.)
@@
Let $G_{1},G_{2},\ldots,G_{t}$ be graphs with $sipco\left(G_{i}\right)\leq k$ for all $i\in[t]$. Then the strong isometric path complexity of $G=\displaystyle\bigoplus\limits_{i=1}^{t}G_{i}$ is at most $3k+18$.
@@
Organization of the paper. In Section 2, we recall some definitions and some results. In Section˜3, we prove Theorems˜2, 6 and 7. In Section˜4, we prove Theorems˜8 and 9. Finally, we conclude in Section˜5.